OpenStudy (anonymous):

There are two numbers of two digits each such that the tens digit of the first number is 3/2 times the tens digit of the second number, while the sum of these two number is 158.find the difference between them?

OpenStudy (anonymous):

X/y=3/2=>x=3 . y=2 x+y=5 x________3______? x+y_____5______158 x=3.158/5=74.8 74.8+y=158 => y=158-74.8 y=123.2

OpenStudy (cwrw238):

let first digit of one number be x then second= 1.5x also 10x + y + 15x + z = 158 25x + y + z = 158

OpenStudy (cwrw238):

can you suggest how we continue? remember that x, y and z must be positive whole numbers

OpenStudy (cwrw238):

its a good bet that x = 6 and y + z = 8

OpenStudy (cwrw238):

in fact that must be the case for if x = 5 then z + y would be 158-125 = 33 which is impossible because x and y are single digits.

OpenStudy (anonymous):

If we assume first no as x and 2nd as y then itbecomes x=3/2y isnt it?

OpenStudy (cwrw238):

yes thats the same as 1.5x

OpenStudy (cwrw238):

no (3/2) x

OpenStudy (cwrw238):

the first digits are 6 and 9 2 numbers are 6y and 9z

OpenStudy (cwrw238):

and y + z = 8 so there is no unique answer to this question possible values of x and y are 3 and 5 ,or 1 and 7 etc

OpenStudy (cwrw238):

- unless i've missed something here....

OpenStudy (anonymous):

Ans is 36 I dnt knw how..

OpenStudy (anonymous):

Can u explain?

OpenStudy (anonymous):

Ans is 36 I dnt knw how..

OpenStudy (anonymous):

If we assume first no as x and 2nd as y then itbecomes x=3/2y isnt it?

OpenStudy (akashdeepdeb):

Is anything else given? ANYTHING?

OpenStudy (anonymous):

Let the 2 numbers be : 10a+b , and 10c+d where a & b are integers from 1 to 9, and c & d from 0 to 9. Then, as given, a=(3/2) c . choice of c is restricted, since 'a' needs to be integer. So, c must be even. 2,4,6,8 options are checked and one can easily arrive at the only possibility of c = 6 and a =9 possible pairs can be (90, 68), (91,67), (92, 66), (93,65), (94,64), (95,63), (96,62), (97,61), (98,60) . Hence, there's no unique answer to the question, but 36 obviously is one of the possibilities : (97,61) So, either you forgot to mention something, or in the options given maybe only 36 is there among the possible answers. I can't help any more in this question.

OpenStudy (akashdeepdeb):

So it is hit and trial! Any other direct way?